OpenStudy (anonymous):

Brothers Bob and Tom Katz buy a tent that has a center pole 6.25 feet high. if the sides of the tent are supposed to make a 50 degree angle with the ground, how wide is the tent?

OpenStudy (nikato):

 have you tried using SOHCAHTOA?

OpenStudy (anonymous):

okay so I let the width of the tent be 2x. tan 50=6.25/x. x=6.25/tan50=5.24. so the tent is 2(5.24) approx. 10.5 ft wide

OpenStudy (nikato):

Yea. That's what I got